OpenStudy (anonymous):

Find the equation of the line that passes through the point (–7, 2) and has a slope m = –5 in point–slope form.

OpenStudy (anonymous):

y-y1=m(x-x1) y-2=-5(x-(-7)) y-2=-5(x+7) y-2=-5x-35 y=-5x-33